OpenStudy (anonymous):

solve the equation: 16x^2+24x+9=-16

OpenStudy (earthcitizen):

i suggest you use quadratic formula here ?

OpenStudy (anonymous):

... ok we have the quadratic equation there: \[\LARGE 16x^2+24x+9+16=0\] \[\LARGE 16x^2+24x+25=0\] now use its formula: \[\LARGE {x_{1/2}} = \frac{{ - b \pm \sqrt {{b^2} - 4ac} }}{{2a}}\] do you know to substitute ? :)

OpenStudy (anonymous):

i got 90.5 but it says that its wrong'

OpenStudy (earthcitizen):

\[x = -3/4-i, -3/4+i\]